Problem (3): In a non-magnetized dielectric the plane wave H field is given by: H(z,t)-2ei(1.5x10%+10%) Amps/m Find an expression for E
Answer #1

The magnetic field is:

\vec{B} = \mu_0 \vec{H} = 2\mu_0 e^{i(1.5\times 10^8t+10^6z)}\hat{x} \enskip T

The electric field amplitude is:

E_0 = cB_0 = 3\times 10^8\times 2\mu_0 = 3\times 10^8\times 2\times 1.257\times 10^{-6} = 754.2 \enskip N/C

The electric field is:

\vec{E}=-754.2e^{(1.5\times 10^8t+10^6z)}\hat{y} \enskip N/C

The direction has been determined using the identity:

\vec{B}=\frac{\hat{k}\times \vec{E}}{c}